# Potential Pitfalls in Visual Models of Tipping Points - And How to Fix Them

### v1

@@ Abstract
Visual models play a crucial role in both science and science communication. However, the distinction between mere analogies and mathematically sound graphical representations is not easy and can be misunderstood not only by laypeople but also within academic literature itself. Moreover, even when the graphical representation exactly corresponds to the mathematical model, its interpretation is often far from obvious. In this paper we discuss the ”potential landscape” visualization commonly used for tipping points in the context of nonlinear dynamics and reveal potential pitfalls, in particular when distinguishing bifurcation induced tipping (B-tipping) from noise-induced tipping (N-tipping). We propose new visualization techniques for tipping dynamics, carefully distinguishing between B- and N-tipping as well as between single systems and ensembles of systems. Explicitly, we apply these visualizations both to molecular cell biology and to climate science in order to reveal the crucial differences in the interpretation of the visual models. We find that it is crucial to explicitly discuss the assumptions made within the visual model and to be aware of the risk of misinterpretation. Based on these findings, we propose as a next step to investigate individual mental models induced by these visualizations in the framework of empirical research.
@@ 1 Introduction
In both science and science communication, visualizations play a crucial role in gaining intuition for a given physical process. Many publications, aimed at both professionals and laypeople, include visualizations to make the abstract and often complex physical systems more accessible and understandable [1, 2, 3, 4, 5, 6, 7, 8, 9]. From a didactic point of view, several aspects have to be taken into account when considering visualizations. First, a distinction must be made between the visualization presented by the author of a given publication with a specific purpose and the individual mental model of the recipient induced by this representation, which may be very different from the author’s intended purpose.
Following Pluta et al. [10], we may adapt his general criteria for ”good models” to the specific application of visualizations. Thus, good visualizations should have high levels of conceptual coherence and clarity, should be compatible with visualizations in other fields, should be appropriately parsimonious, and should be consistent with empirical evidence.
@@
It is important to note that the recipient’s mental model depends not only on the visualization, but also on their prior knowledge. The Students’ Understanding of Models in Science instrument (SUMS) was developed in [11] by Treagust et al. to gain insight into these mental models. Using exploratory factor analysis, Treagust et al. identified five factors, namely models as ”multiple representations”, as ”exact replicas” and as ”explanatory tools”, the ”uses of scientific model”, and the ”changing nature of models”. Later, Ubben and Heusler [12] combined these instruments with context-specific questions related to models of the atomic shell.
Using exploratory factor analysis, two main factors were identified with regard to this context:
@@
The first factor is Fidelity of Gestalt (FG), which describes the extent to which models are understood to be accurate visual representations of phenomena or accurate representations of how things look, and the extent to which the gestalt of the models is perceived to be accurate.
The second factor is Functional Fidelity (FF), which describes the extent to which models are understood as appropriate descriptions of how phenomena work, and the extent to which the underlying abstract functionality of the models is perceived as accurate. In subsequent studies, these two independent dimensions FF and FG have been empirically confirmed in various areas of physics [13, 14] and can be assumed to apply rather generally.
@@
In this paper we focus our discussion on nonlinear dynamics for several reasons. Applications of nonlinear dynamics range from climate science to cell dynamics to models of brain activity and artificial intelligence, making it one of the most important frameworks in contemporary physics. For this reason, the communication of scientific results to the general public is particularly important in this context, and indeed visualizations are a crucial communication technique. One of the most popular visualization techniques used to convey the key ideas of nonlinear dynamics is a ”potential landscape” (see Fig.2). The ”point particle” (which represents the state of the system) moves towards a ”stable” minimum within the potential. While the strong point of this visualization is its universality, there is a price to pay for that: In each application at hand, the interpretation of the visualization is rather different. What is the meaning of the ”potential”? How should we interpreted ”the point particle”? What mechanisms may change the behaviour of the ball, in particular, the transition from one minimum of the potential to another (also called ”tipping behaviour”)?
@@
Particularly in the context of climate science, the topic of ”tipping points” attracts much attention [15, 16]. Climate science seeks to understand the dynamics of the components of the Earth’s climate system and to model and predict future changes in the Earth’s climate [17]. These climate components (e.g. atmosphere, AMOC, Greenland Ice, Monsoons, Amazon rainforest) are highly complex systems that also exhibit a wide range of nonlinear properties, such as stochastic and chaotic behavior, hysteresis and tipping points. Therefore, visualization techniques are ubiquitous to simplify the complex behavior of climate systems and to help understand and communicate findings effectively. For example, many major publications on climate tipping points also make use of potential landscape visualizations to explain system dynamics [18, 16, 19, 1, 3, 20].
However, the same visualization ideas can also be applied to topics such as the transition of a cell from a pluripotent stem cell to different possible somatic cell states (called differentiation) in the field of molecular cell biology, where nonlinear dynamics also play an essential role [21, 4]. Many publications in this field also rely on the use of potential landscape visualizations to explain and communicate these differentiation and cell fate decision processes [21, 22, 4, 5, 6, 23, 24].
@@
We choose to focus on these two particular fields of science because on the one hand both exhibit similar stochastic nonlinear behaviour, in particular noise- and bifurcation-induced tipping. On the other hand, there are subtle differences in the interpretation of their respective visualizations that we want to highlight (see Sect. 4). As we will show later in this paper, visualizations of these systems can sometimes be ambiguous, misleading, and even problematic for various reasons. We will identify several misleading and problematic visualizations, either in their communicative purpose intended by the author, or in possible misinterpretations in terms of individual mental models, or even both. We then propose new visualization techniques for nonlinear stochastic processes from the fields of molecular cell biology and climate science as well as point out subtle difficulties in their interpretation. For that purpose, we carefully distinguish visualizations along two dimensions of interpretation: (i) interpretation as ensemble/individual system (ii) interpretation as deterministic/stochastic system (see Fig. 1). While we apply our visualization to the context of climate science and cell dynamics, we argue that our visualization scheme is rather universal and may be adopted to any other context within the field of nonlinear dynamics. Further empirical research will be necessary to elucidate the mental models induced by these new visualizations.
@@
This paper is organised as follows: First, we discuss analogies between classical mechanics and nonlinear dynamics, potential landscapes as a shared visualization technique and limitations of their analogies. We then give a brief introduction to the mathematical background of deterministic and stochastic nonlinear dynamics and in particular tipping behavior. More specifically, we distinguish between deterministic and stochastic systems as well as single and ensemble systems - and the resulting consequences for visualizations and their interpretation. Next, we focus on tipping behavior in climate systems and cell differentiation processes and potential problems of their visualizations. We then propose our own visualizations, carefully distinguishing between deterministic and stochastic as well as single and ensemble systems. Finally, we suggest further empirical research to better understand the mental models these visualizations induce, and emphasize the need for clear and accurate visual representations in scientific communication.
@@ 2 The Analogy Between Classical Mechanics and Nonlinear Dynamics
Much of the intuition about nonlinear dynamics is based on analogies to classical mechanics (see e.g. Strogatz [25]). In this context, consider the visualizations shown in Fig. 2, which can be found in similar form in popular and academic literature.
Figure 2: (a) Analogy from classical mechanics of a one-dimensional bistable system with valleys (stable equilibria) at points $A$ and $B$ separated by a saddle (instable equilibrium) at point $S$. (b) Abstract functional potential of the same shape with the same equilibria.
The ”potential landscape” shown here describes the (potential) energy of a system as a function of one variable with the position of the system state, represented as spheres within the landscape, corresponding to the current energy of the system. If the potential landscape is understood by the recipient not as an abstract functional object (functional fidelity), but as a real shape of a mountain landscape (fidelity of gestalt), misinterpretations can arise. Taken literally, one could assume that frictionless motion in a quadratic ”potential landscape” $z(x)=h(x)=ax^{2}$ leads to harmonic oscillation, which is not the case. The functional relation $z=h(x)$ leads to a constraint in the equation of motion, which can be treated using Lagrange parameters. The force $F_{T}=mg\sin(\phi)=mgh^{\prime}(\sqrt{1+(h^{\prime})^{2}})$ with $h^{\prime}=dh/dx=\tan(\phi)$ would then be nonlinear. On the other hand, if we took ”fidelity of gestalt” literally, we could determine the shape $h(x)$ of a real ”mountain landscape” that would indeed lead to a harmonic oscillation. For this we must require that the force $mg\sin(\phi)$ is proportional to the path length $s=\int\sqrt{dx^{2}+dz^{2}}$. This results in the following functional equation for $z=h(x)$:
[…]
with some proportionality constant $D$. While this is an amusing and perhaps unexpected problem of classical mechanics, it is obvious that the naive (FG)-interpretation of the potential as some real ”mountain landscape” is a dead end. Instead, this analogy should be understood only in an abstract, functional sense (FF).
[…]
@@
Again by analogy, the concept of ”potential” is further generalized when we consider models in nonlinear dynamics. Now the ”state” of the system is no longer a point particle like in classical mechanics, but can have a vast number of interpretations - e.g. as a particular state of cell dynamics, brain activity, strength of the AMOC, percentage of forest cover of the Amazon rainforest, etc. For deterministic processes of the type
@@
the interpretation of a ”particle” (a system) at position ${\bm{x}}$ (a specific system state) in a ”potential” $V({\bm{x}},t,\eta)$ describing the dynamics of the system state for a given set $\eta$ of external parameters can be visualized as shown in Fig 2 (b). At any given time, the state is described by the vector ${\bm{x}}$ describing the unique and deterministic system state driven by the dynamics. Fixed points of the dynamics, where $\dot{\bm{x}}=f({\bm{x}},t,\eta)=0$, are represented by extremal points of the potential, with stable fixed points corresponding to minima.
Remarkably, some general mechanisms can be derived that underlie all these different types of dynamics, revealing a certain universality of nonlinear dynamics. From this point of view, it is clear that modeling in terms of ”potentials” must be seen as a rather abstract concept.
@@ 3 Mathematical Background of Nonlinear Dynamics and Tipping Behavior
In bi- or multistable systems, i.e. systems that have more than one stable state, B-tipping can occur when one of the stable states disappears, changes its stability, or the system is forced into a qualitatively different stable state due to a local bifurcation. Bifurcations are caused by internal or external parameter changes corresponding to a change of the physical conditions of the system or its environment. In terms of the potential landscape, let $\bm{x}$ be the system state and $\eta$ an external forcing parameter. The values $(\bm{x_{b}},\eta_{b})$ at which the bifurcation occurs are called bifurcation points. Whithin the potential landscape, a stable minimum can then become unstable once this bifurcation points is reached and surpassed.
In contrast, N-tipping can occur when the current attractor of a system decreases in stability. The system is then more susceptible to small perturbations caused by stochastic fluctuations, i.e. noise. These can cause the system to leave the neighborhood of its stable attractor and cross the stability threshold, resulting in a transition to another stable state, even if the potential itself does not change its shape. N-tipping is caused by stochastic forcing, which drives the system state from the neighborhood of a stable fixed point across a stability boundary into the basin of attraction of another stable fixed point. To model the stochastic forcing of a dynamical system, an additive, state-independent noise term can be included in the model, resulting in the stochastic differential equation (SDE) [29]
@@
Here, the added term describes the stochastic forcing, where $\bm{\sigma}$ is the noise amplitude matrix and $d\bm{W}_{t}$ is the multidimensional increment of a Wiener process, i.e., white noise. Since the system is now non-deterministic, instead of individual system trajectories, the probability density function $p(\bm{x},\eta,t)$ of the system states must be computed. This can be done by solving the Fokker-Planck equation [29]
$\frac{\partial p(\bm{x},\eta,t)}{\partial t}=-\sum_{i=1}^{N}{\frac{\partial}{\partial x_{i}}}\bigl{[}f_{i}(\bm{x},\eta,t)\,p(\bm{x},\eta,t)\bigr{]}+\sum_{i=1}^{N}\sum_{j=1}^{N}{\frac{\partial^{2}}{\partial x_{i}\,\partial x_{j}}}\bigl{[}D_{ij}\,p(\bm{x},\eta,t)\bigr{]}$ (5)
with the diffusion Matrix $\bm{D}=\frac{1}{2}\bm{\sigma\sigma}^{\text{T}}$.
One of the crucial questions we want to address in this paper is: How can stochastic behaviour be incorporated into the visualization of the potential landscape? As a preparation to this question, we again recall the original analogy to classical mechanics. In many-particle physics and in thermodynamics, even if the underlying classical dynamics is deterministic, it is natural to introduce statistical methods to evaluate mean values and variations. The concept of a phase space density $\rho({\bm{x}},{\bm{p}},t)$ is central to this statistical description of time dynamics. Adding noise to a single-particle system can effectively describe fluctuations due to random interactions in many-particle systems, and we can therefore reduce its many-particle interaction to a model of a single particle with noise. By analogy, we can also add noise to nonlinear systems as an effective description of unknown interactions with the environment. Now the state of the system can only be described by a probability density $p({\bm{x}},t)$, like in a diffusive process. For this reason, the interpretation of the visualization of the system state as a single ”point particle” (as shown in Fig. 2) becomes undefined for N-tipping, and indeed this visualization is imprecise since it does not show the stochastic nature of the process.
Dynamical systems in the real world are almost always subject to internal or external forcing, leading to possible interactions between B- and N-tipping. Small fluctuations of the system state will not lead to tipping as long as the shape of the potential landscape does not change and the potential well is much higher than the magnitude of the fluctuation. However, when the system approaches a bifurcation point caused by a change in the external parameters $\eta$, the system loses resilience (i.e. the potential well flattens) and small fluctuations can then cause tipping even before the bifurcation point is reached. In the case of strong forcing and large fluctuations, tipping becomes possible even far away from a bifurcation point.
@@
In this section, we propose our own visualization technique, carefully distinguishing ensemble from single systems and stochastic from deterministic behaviour, as shown in Fig. 1. Later on, we apply these visualizations to cell dynamics and to tipping points in the climate system.
First, we extend the simple example of Brownian motion to a more complex nonlinear system with tipping behavior. Here, the same distinction between individual systems and ensembles as well as their interpretations must be made.
In the deterministic case without stochastic forcing, the ensemble behaves identically to a single system state, with each member trivially following exactly the same path determined by the underlying deterministic system dynamics. These cases are often visualized by a potential landscape in which the system state is represented by a sphere. B-tipping then leads to a change in stability, e.g. the flattening of a potential well, as visualized in Figure 4.
@@
With the introduction of stochastic forcing, an additional diffusion of system states (for an ensemble) or the probability density (for a single system) emerges.
In Fig. 5, we present a potential landscape visualization of a nonlinear dynamical process under stochastic forcing (B- and N-tipping) for an ensemble of systems (a), (b) and a single system (c), (d).
Figure 5: (a), (b): Ensemble interpretation of a stochastic nonlinear process. (c), (d): Probabilistic interpretation of the same dynamics. (a), (c): Situation far from the bifurcation. (b), (d): Situation near the bifurcation. Due to the stochastic forcing, there is a certain probability of N-tipping even before the B-tipping point is reached.
In an ensemble under stochastic forcing, different individual system states diffuse to different fixed points of the dynamics, as shown in Fig. 5(a), (b). In this interpretation, each member of the ensemble can be said to behave differently, and we can characterize this behavior in the limit of large $N$ by a cloud of ”particles”, each representing an individual system. Fig. 5 also shows the state density of the ensemble (which coincides with the probability density of the individual system in the limit of large ensemble size $N$). In this paper, as a prototypical example of the applicability of the ensemble interpretation, we discuss molecular cell dynamics, where each ”particle” represents the state of a single cell.
In contrast, Fig. 5(c), (d) show the diffusion of the probability density within the potential landscape for a $single$ system. Here, since $N=1$, the ensemble interpretation is no longer possible. Instead, the probability density function that describes the state probability of a single system must be used. Our representation shows the probability distribution within the potential landscape. In this paper, as a prototypical example of the applicability of the stochastic interpretation of a single classical system, we consider a model of one component of the climate system, the Amazon rainforest.

### v2

@@ Abstract
Visual models play a crucial role in both science and science communication. However, the distinction between mere analogies and mathematically sound graphical representations is not easy and can be misunderstood not only by laypeople but also within academic literature itself. Moreover, even when the graphical representation exactly corresponds to the mathematical model, its interpretation is often far from obvious. In this paper we discuss the potential landscape visualization commonly used for tipping points in the context of nonlinear dynamics and reveal potential pitfalls, in particular when distinguishing bifurcation induced tipping (B-tipping) from noise-induced tipping (N-tipping).
We propose new visualization techniques for tipping dynamics, carefully distinguishing between B- and N-tipping as well as between single systems and ensembles of systems. Explicitly, we apply these visualizations both to molecular cell biology and to climate science in order to reveal the crucial differences in the interpretation of the visual models. We find that it is crucial to explicitly discuss the assumptions made within the visual model and to be aware of the risk of misinterpretation. These findings apply to a wide range of readership, from graduate students - as some general knowledge of nonlinear systems is required - to research professionals working in the field of nonlinear sciences. This paper provides the theoretical groundwork for these new visualizations. As a next step, we propose to investigate the individual mental models that might be induced by these visualizations using empirical research that builds upon these findings.
@@ 1 Introduction
In both science and science communication, visualizations play a crucial role in gaining intuition for a given physical process. Many publications, aimed at both professionals and laypeople, include visualizations to make the abstract and often complex physical systems more accessible and understandable [37, 1, 18, 4, 11, 26, 24, 5, 10]. From a didactic point of view, several aspects have to be taken into account when considering visualizations. First, a distinction must be made between the visualization presented by the author of a given publication with a specific purpose and the individual mental model the recipient might form based on this representation, which may be very different from the author’s intended purpose.
@@
To analyze the potential pitfalls of visualizations, it is helpful to use a theoretical framework that distinguishes between different aspects of a viewer’s interpretation. Pluta et al. [28] provide general criteria for ”good models,” which we can adapt to visualizations: they should have high levels of conceptual coherence and clarity, should be compatible with visualizations in other fields, should be appropriately parsimonious, and should be consistent with empirical evidence. It is important to note that the recipient’s mental model can not only depend on the visualization, but also on their prior knowledge. The Students’ Understanding of Models in Science instrument (SUMS) was developed by Treagust et al. [12] to gain insight into these mental models. Using exploratory factor analysis, Treagust et al. identified five factors, namely models as ”multiple representations”, as ”exact replicas” and as ”explanatory tools”, the ”uses of scientific model”, and the ”changing nature of models”. Later, Ubben and Heusler [30] combined these instruments with context-specific questions related to models of the atomic shell.
@@
Using exploratory factor analysis, two main factors were identified with regard to this context: The first factor is Fidelity of Gestalt (FG), which describes the extent to which models are understood to be accurate visual representations of phenomena or accurate representations of how things look, and the extent to which the gestalt of the models is perceived to be accurate. The second factor is Functional Fidelity (FF), which describes the extent to which models are understood as appropriate descriptions of how phenomena work, and the extent to which the underlying abstract functionality of the models is perceived as accurate.
This FF/FG distinction is crucial because many pitfalls in scientific visuals arise from a ”category error,” where a viewer interprets a functional model with high fidelity of gestalt. This framework, which has been empirically confirmed in various areas of physics [31, 32], will guide our critique of existing visualizations and our proposal for new ones. It also provides the necessary vocabulary for the future empirical research we advocate for in our conclusion.
@@
In this paper we focus our discussion on nonlinear dynamics for several reasons. Applications of nonlinear dynamics range from climate science to cell dynamics to models of brain activity and artificial intelligence, making it one of the most important frameworks in contemporary physics. For this reason, the communication of scientific results to the general public is particularly important in this context. The inherent difficulty of conveying counter-intuitive concepts such as ”bifurcations”, ”stochasticity”, and ”tipping points” with words alone makes visualizations an essential, yet challenging, communication technique. One of the most popular visualization techniques used to convey the key ideas of nonlinear dynamics is a ”potential landscape” (see Fig.2). The ”point particle” (which represents the state of the system) moves towards a ”stable” minimum within the potential. While the strong point of this visualization is its universality, there is a price to pay for that: In each application at hand, the interpretation of the visualization is rather different. What is the meaning of the ”potential”? How should we interpret ”the point particle”? What mechanisms may change the behavior of the ball, in particular, the transition from one minimum of the potential to another (also called ”tipping behavior”)?
@@
Particularly in the context of climate science, the topic of ”tipping points” attracts much attention [19, 20]. Climate science seeks to understand the dynamics of the components of the Earth’s climate system and to model and predict future changes in the Earth’s climate [3]. These climate components (e.g., atmosphere, AMOC, Greenland Ice, Monsoons, Amazon rainforest) are highly complex systems that also exhibit a wide range of nonlinear properties, such as stochastic and chaotic behavior, hysteresis and tipping points. Therefore, visualization techniques are ubiquitous to simplify the complex behavior of climate systems and to help understand and communicate findings effectively. For example, many major publications on climate tipping points also use potential landscape visualizations to explain system dynamics [35, 20, 21, 37, 18, 23]. However, the same visualization ideas can also be applied to topics such as the transition of a cell from a pluripotent stem cell to different possible somatic cell states (called differentiation) in the field of molecular cell biology, where nonlinear dynamics also play an essential role [39, 4]. Many publications in this field also rely on the use of potential landscape visualizations to explain and communicate these differentiation and cell fate decision processes [39, 40, 4, 11, 26, 6, 7].
@@
We choose to focus on these two particular fields of science because on the one hand, both exhibit similar stochastic nonlinear behavior, in particular noise- and bifurcation-induced tipping. On the other hand, there are subtle differences in the interpretation of their respective visualizations that we want to highlight (see Sec. 4). As we will show later in this paper, visualizations of these systems can sometimes be ambiguous, misleading, and even problematic for various reasons. We will identify several misleading and problematic visualizations, either in their communicative purpose intended by the author, or in possible misinterpretations in terms of individual mental models, or even both. We then propose new visualization techniques for nonlinear stochastic processes from the fields of molecular cell biology and climate science as well as point out subtle difficulties in their interpretation. For that purpose, we carefully distinguish visualizations along two dimensions of interpretation: (i) interpretation as ensemble/individual system (ii) interpretation as deterministic/stochastic system (see Fig. 1). While we apply our visualization to the context of climate science and cell dynamics, we argue that our visualization scheme is generalizable and may be adopted to any other context within the field of nonlinear dynamics. Further empirical research will be necessary to elucidate the mental models induced by these new visualizations.
@@
This paper is organized as follows: First, we discuss analogies between classical mechanics and nonlinear dynamics, potential landscapes as a shared visualization technique and limitations of their analogies. We then give a brief introduction to the mathematical background of deterministic and stochastic nonlinear dynamics and in particular tipping behavior. More specifically, we distinguish between deterministic and stochastic systems as well as single and ensemble systems - and the resulting consequences for visualizations and their interpretation. Next, we focus on tipping behavior in climate systems and cell differentiation processes and potential problems of their visualizations. We then propose our own visualizations, carefully distinguishing between deterministic and stochastic as well as single and ensemble systems. Finally, we suggest further empirical research to better understand the mental models these visualizations might induce, and emphasize the need for clear and accurate visual representations in scientific communication.
@@ 2 The Analogy Between Classical Mechanics and Nonlinear Dynamics
Much of the intuition about nonlinear dynamics is based on analogies to classical mechanics (see e.g., Strogatz [38]). In this context, consider the visualizations shown in Fig. 2, which can be found in similar form in popular and academic literature.
Figure 2: (a) Analogy from classical mechanics of a one-dimensional bistable system with valleys (stable equilibria) at points AA and $B$ separated by a saddle (instable equilibrium) at point $S$. (b) Abstract functional potential of the same shape with the same equilibria. Visualizations of this type are a common introductory metaphor for bistability and are found in many standard texts on nonlinear dynamics (e.g., Strogatz [38])
The ”potential landscape” shown here describes the (potential) energy of a system as a function of one variable with the position of the system state, represented as spheres within the landscape, corresponding to the current energy of the system. If the potential landscape is understood by the recipient not as an abstract functional object (functional fidelity), but as a real shape of a mountain landscape (fidelity of gestalt), misinterpretations can arise. For instance, while frictionless motion in an abstract quadratic potential energy well $V(x)\propto x^{2}$ indeed results in simple harmonic oscillation for any amplitude, this is not true if one misinterprets the visualization as a real physical track with a quadratic shape $h(x)=ax^{2}$. In the latter case, the motion is constrained to the curve and the restoring force along the path becomes a nonlinear function of position, which deviates from simple harmonic oscillation for large amplitudes. The constraint in the equation of motion introduced by the functional relation $z=h(x)$, can be treated using Lagrange parameters. The force
$F_{T}=mg\sin(\phi)=mg\frac{h^{\prime}}{\sqrt{1+(h^{\prime})^{2}}}\quad\text{with}\quad h^{\prime}=dh/dx=\tan(\phi)$ (1)
would then be nonlinear. On the other hand, if we took ”fidelity of gestalt” literally, we could determine the shape $h(x)$ of a real ”mountain landscape” that would indeed lead to a harmonic oscillation. For this we must require that the force $mg\sin(\phi)$ is proportional to the path length $s=\int\sqrt{dx^{2}+dz^{2}}$. This results in the following functional equation for $z=h(x)$:
$\frac{h^{\prime}}{\sqrt{1+(h^{\prime})^{2}}}=D\int_{0}^{s}\sqrt{1+(h^{\prime})^{2}}dx$ (2)
with some proportionality constant $D$. While this is an amusing and perhaps unexpected problem of classical mechanics, it is obvious that the naive (FG)-interpretation of the potential as some real ”mountain landscape” is a dead end, as the visual shape of a landscape and its abstract physical function are not the same. Instead, this analogy should be understood only in an abstract, functional sense (FF).
Even on this level, there are various interpretations of the ”potential”: In Newtonian mechanics, the potential energy (with the physical unit ”Joule”) is related to the physical concept of ”work” in a conservative force field $\bm{F}$ via $V(\bm{x})=-\int^{\bm{x}}{\bm{F}}d{\bm{s}}$, equivalently,
@@
Again by analogy, the concept of ”potential” is further generalized when we consider models in nonlinear dynamics. Now the ”state” of the system is no longer a point particle like in classical mechanics, but can have a vast number of interpretations - for example as a particular state of cell dynamics, brain activity, strength of the AMOC, percentage of forest cover of the Amazon rainforest, etc. For deterministic processes of the type
@@
the interpretation of a ”particle” (a system) at position $\bm{x}$} (a specific system state) in a ”potential” $V({\bm{x}},t,\eta)$ describing the dynamics of the system state for a given set $\eta$ of external parameters can be visualized as shown in Fig 2 (b). At any given time, the state is described by the vector $\bm{x}$} describing the unique and deterministic system state driven by the dynamics. Fixed points of the dynamics, where $\dot{\bm{x}}=f({\bm{x}},t,\eta)=0$, are represented by extremal points of the potential, with stable fixed points corresponding to minima. Remarkably, some general mechanisms can be derived that underlie all these different types of dynamics, revealing a certain universality of nonlinear dynamics. From this point of view, it is clear that modeling in terms of ”potentials” must be seen as a rather abstract concept.
@@ 3 Mathematical Background of Nonlinear Dynamics and Tipping Behavior
In bi- or multistable systems, i.e., systems that have more than one stable state, B-tipping can occur when one of the stable states disappears, changes its stability, or the system is forced into a qualitatively different stable state due to a local bifurcation. Bifurcations are caused by internal or external parameter changes corresponding to a change of the physical conditions of the system or its environment. In terms of the potential landscape, let $\bm{x}$ be the system state and $\eta$ an external forcing parameter. The values $(\bm{x_{b}},\eta_{b})$ at which the bifurcation occurs are called bifurcation points. Within the potential landscape, a stable minimum can then become unstable once this bifurcation points is reached and surpassed.
In contrast, N-tipping can occur when the current attractor of a system decreases in stability. The system is then more susceptible to small perturbations caused by stochastic fluctuations, i.e., noise. These can cause the system to leave the neighborhood of its stable attractor and cross the stability threshold, resulting in a transition to another stable state, even if the potential itself does not change its shape. N-tipping is caused by stochastic forcing, which drives the system state from the neighborhood of a stable fixed point across a stability boundary into the basin of attraction of another stable fixed point. To model the stochastic forcing of a dynamical system, an additive, state-independent noise term can be included in the model, resulting in the stochastic differential equation (SDE) [16]
@@
Here, the added term describes the stochastic forcing, where $\bm{\sigma}$ is the noise amplitude matrix and $d\bm{W}_{t}$ is the multidimensional increment of a Wiener process, i.e., white noise. Since the system is now non-deterministic, individual system trajectories are no longer sufficient. Instead, we must consider the evolution of the probability density function $p(\bm{x},\eta,t)$. The mathematical tool that governs the dynamics of this probability density is the Fokker-Planck equation [16]. While the full formalism is complex, we present it here briefly to mathematically ground our subsequent discussion of stochastic visualizations. The Fokker-Planck equation with the diffusion Matrix $\bm{D}=\frac{1}{2}\bm{\sigma\sigma}^{\text{T}}$ is
$\frac{\partial p(\bm{x},\eta,t)}{\partial t}=-\sum_{i=1}^{N}{\frac{\partial}{\partial x_{i}}}\bigl{[}f_{i}(\bm{x},\eta,t)\,p(\bm{x},\eta,t)\bigr{]}+\sum_{i=1}^{N}\sum_{j=1}^{N}{\frac{\partial^{2}}{\partial x_{i}\,\partial x_{j}}}\bigl{[}D_{ij}\,p(\bm{x},\eta,t)\bigr{]}.$ (6)
One of the crucial questions we want to address in this paper is: How can stochastic behavior be incorporated into the visualization of the potential landscape? As a preparation to this question, we again recall the original analogy to classical mechanics. In many-particle physics and in thermodynamics, even if the underlying classical dynamics is deterministic, it is natural to introduce statistical methods to evaluate mean values and variations. The concept of a phase space density $\rho({\bm{x}},{\bm{p}},t)$ is central to this statistical description of time dynamics. Adding noise to a single-particle system can effectively describe fluctuations due to random interactions in many-particle systems, and we can therefore reduce its many-particle interaction to a model of a single particle with noise. By analogy, we can also add noise to nonlinear systems as an effective description of unknown interactions with the environment. Now the state of the system can only be described by a probability density $p({\bm{x}},t)$, like in a diffusive process. For this reason, the interpretation of the visualization of the system state as a single ”point particle” (as shown in Fig. 2) becomes undefined for N-tipping, and indeed this visualization is imprecise since it does not show the stochastic nature of the process.
Dynamical systems in the real world are almost always subject to internal or external forcing, leading to possible interactions between B- and N-tipping. Small fluctuations of the system state will not lead to tipping as long as the shape of the potential landscape does not change and the potential well is much higher than the magnitude of the fluctuation. However, when the system approaches a bifurcation point caused by a change in the external parameters $\eta$, the system loses resilience (i.e., the potential well flattens) and small fluctuations can then cause tipping even before the bifurcation point is reached. In the case of strong forcing and large fluctuations, tipping becomes possible even far away from a bifurcation point.
@@
In this section, we propose our own visualization technique, carefully distinguishing ensemble from single systems and stochastic from deterministic behavior, as shown in Fig. 1. Later on, we apply these visualizations to cell dynamics and to tipping points in the climate system. First, we extend the simple example of Brownian motion to a more complex nonlinear system with tipping behavior. Here, the same distinction between individual systems and ensembles as well as their interpretations must be made.
In the deterministic case without stochastic forcing, the ensemble behaves identically to a single system state, with each member trivially following exactly the same path determined by the underlying deterministic system dynamics. These cases are often visualized by a potential landscape in which the system state is represented by a sphere. B-tipping then leads to a change in stability, e.g., the flattening of a potential well, as visualized in Figure 4.
@@
With the introduction of stochastic forcing, an additional diffusion of system states (for an ensemble) or the probability density (for a single system) emerges. In Fig. 5, we present a potential landscape visualization of a nonlinear dynamical process under stochastic forcing (B- and N-tipping) for an ensemble of systems (a, b) and a single system (c, d).
Figure 5: (a, b): Ensemble interpretation of a stochastic nonlinear process. (c, d): Probabilistic interpretation of the same dynamics. (a, c): Situation far from the bifurcation. (b, d): Situation near the bifurcation. Due to the stochastic forcing, there is a certain probability of N-tipping even before the B-tipping point is reached.
It is important to note that these proposed visualizations are designed specifically to address the challenge of representing stochasticity and distinguishing between single-system and ensemble interpretations. While the underlying ”potential landscape” metaphor is retained to provide a familiar visual scaffold, the key innovation lies in the representation of the system state itself. By visually differentiating between the cloud of points representing an ensemble’s distribution of states (Fig. 5a, b) and the probability density shading representing a single stochastic system (Fig. 5c, d), our approach makes these crucial, and often overlooked, distinctions explicit. This is a critical improvement, as traditional visualizations often use a single, ambiguous sphere to represent the system state, regardless of whether it is a deterministic or stochastic system, a single system or an entire ensemble. Differentiating these cases is essential for accurately depicting the interplay between bifurcation-induced and noise-induced tipping.
In an ensemble under stochastic forcing, different individual system states diffuse to different fixed points of the dynamics, as shown in Fig. 5a, b. In this interpretation, each member of the ensemble can be said to behave differently, and we can characterize this behavior in the limit of large $N$ by a cloud of ”particles”, each representing an individual system. Fig. 5 also shows the state density of the ensemble (which coincides with the probability density of the individual system in the limit of large ensemble size $N$). In this paper, as a prototypical example of the applicability of the ensemble interpretation, we discuss molecular cell dynamics, where each ”particle” represents the state of a single cell.
In contrast, Fig. 5c, d show the diffusion of the probability density within the potential landscape for a singlesingle system. Here, since $N=1$, the ensemble interpretation is no longer possible. Instead, the probability density function that describes the state probability of a single system must be used. Our representation shows the probability distribution within the potential landscape. In this paper, as a prototypical example of the applicability of the stochastic interpretation of a single classical system, we consider a model of one component of the climate system, the Amazon rainforest.